OpenStudy (anonymous):

An office supply store sells wide-ruled notebook paper for $1.75 per pound. College-rule notebook paper costs $2.20 per pound. How many pounds of wide-ruled notebook paper are needed to make a 5-pound combination of paper that costs $2.11?

OpenStudy (anonymous):

@Arfney

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

x = wide ruled paper = 1.75 y = college ruled paper = 2.20 wide rulled paper+ college rulled paper = 5 So we have the equations 1.75x + 2.20y = 2.11 and x + y = 5

OpenStudy (anonymous):

So if we make everything a whole number by multiplying by 100 we get 175x +220y = 211 then using elimination (x + y = 5) *175 175x +220y = 211 175x + 175y = 875 subtracting them we get 45y = -6.64 dividing by 45 we get y = .14755 then plug that into the second equation x + y = 5 x + .14755 = 5 x = 4.8524 I think I may have done something wrong...

OpenStudy (anonymous):

Yeah.. I'm not really sure here.....

OpenStudy (anonymous):

I think the answer is x = 19.755 and y = -14.755

OpenStudy (anonymous):

I did it on paper and that is what I got.

OpenStudy (anonymous):

okay, thanks!

OpenStudy (anonymous):

Also to my answer that I posted 45y = -6.64 <-------- should be -664 not -6.64 So then dividing by 45 we get y = 14.755 then plug that into the second equation x + y = 5 x - 14.755 = 5 x = 19.8524 So that is right!